# Flow-Based Conformal Predictive Distributions

### v1

@@ Abstract
Conformal prediction provides a distribution-free framework for uncertainty quantification via prediction sets with exact finite-sample coverage. In low dimensions these sets are easy to interpret, but in high-dimensional or structured output spaces they are difficult to represent and use, which can limit their ability to integrate with downstream tasks such as sampling and probabilistic forecasting. We show that any differentiable nonconformity score induces a deterministic flow on the output space whose trajectories converge to the boundary of the corresponding conformal prediction set. This leads to a computationally efficient, training-free method for sampling conformal boundaries in arbitrary dimensions. Boundary samples can be reconformalized to form pointwise prediction sets with controlled risk, and mixing across confidence levels yields conformal predictive distributions whose quantile regions coincide exactly with conformal prediction sets. We evaluate the approach on PDE inverse problems, precipitation downscaling, climate model debiasing, and hurricane trajectory forecasting.
@@ 1 Introduction
In this work, we propose a constructive alternative based on mathematical flows [36]. We show that, for any differentiable score, we can define a deterministic dynamical system on $\mathcal{Y}$ whose trajectories globally attract to the target level set. The induced flow, called a nonconformity flow, is uniquely determined by the gradient of the nonconformity score, and has two key properties. First, it requires no training, auxiliary modeling, or transport map estimation and can be computed directly from the score. Second, for almost any initialization, trajectories converge exponentially fast to the boundary of the target conformal prediction set. Nonconformity flows yield an exact, computationally efficient mechanism for sampling conformal prediction set boundaries in arbitrary dimensions.
These samples can be used directly to construct risk-controlling prediction bands [5] and alternative UQ representations. Furthermore, by mixing across confidence levels we can naturally induce a class of admissible predictive distributions called conformal predictive distributions (CPDs) for which every conformal prediction set $C_{\alpha}(x)$ is an exact quantile region. Constructed in this way, CPDs are a highly scalable Monte Carlo alternative to analytical conformal predictive systems [42] and optimal transport maps [32] for lifting conformal prediction from sets to predictive distributions.
We evaluate nonconformity flows and conformal predictive distributions on a range of structured regression tasks, including elliptic PDE inverse problems, precipitation downscaling, climate model debiasing, and trajectory forecasting (Section 4). We show that nonconformity flows quickly and reliably converge across a range of data modalities and score functions and that the derived risk-controlled bands are tight and adaptive. CPDs are shown to be competitive with baseline uncertainty quantification (UQ) methods and can act as fully controllable rare-event simulators (Figure 6).
@@ 3 Method
Section 3 develops our method in three stages. First, we introduce nonconformity flows to efficiently sample from $\partial C_{\alpha}(x)=\{y\in\mathcal{Y}:S(x,y)=\tau_{\alpha}\}$ in arbitrary dimensions given any differentiable score (Section 3.1). We then show how to convert these into risk-controlling prediction bands (Section 3.2). Finally, we show how mixing across confidence levels induces a class of calibrated conformal predictive distributions (CPDs), allowing us to transform conformal sets in to fully fledged predictive distributions (Section 3.3).
@@ 3.1 Sampling conformal boundaries
$T_{\varepsilon}(y_{0})=\frac{1}{\lambda}\log\frac{|S(y(0))-\tau_{\alpha}|}{\varepsilon}.$
@@
3.2 Risk–controlling Prediction Bands
$P(\ell_{\eta}(X,Y)\leq\delta\bigr)\geq 1-\alpha.$ (9)
@@ 4 Numerical Experiments
We evaluate the proposed framework along three axes corresponding to the theoretical developments in Section 3: (i) convergence and robustness of the nonconformity flow (Section 4.1), (ii) construction of risk-controlling prediction bands via reconformalization (Section 4.1), and (iii) conformal predictive distributions (CPDs) as calibrated predictive objects (Section 4.2). Unless otherwise stated, definitions of score functions, data generation, preprocessing, model architectures, and evaluation metrics are deferred to Appendices D.1 - D.6.
@@ 4.1 Evaluating Conformal Level Sets
We first evaluate whether the proposed flow (Equation 5) converges to within tolerance ($\varepsilon=10^{-6}$) with $\lambda$ set per trajectory according to the $\varepsilon$-hitting time equation (Equation 3.2). We consider isotropic Gaussian, anisotropic Student-t ($df=3$), 2D Gaussian processes, and 2D GP upsampling targets across a range of dimensions 51005-100 (isotropic Gaussians, anisotropic Student-t) or intrinsic dimension range $\ell=0.01-0.2$ (2D GP tasks) (Appendix D.1). For the vector regression tasks, we train a multilayer perceptron (Appendix D.2) and for the operator tasks, a 2D Fourier Neural Operator (Appendix D.2). All models trained on a single A100 GPU for 25-100 epochs until convergence using the Adam optimizer with learning rate $\gamma=10^{-3}$.
@@
Figure 3a, shows the log score error after 20 integration steps broken out by score function (Appendix D.3D.4) with the black line indicating the target tolerance. Across all settings and tasks the conformal flow sampler converges to within tolerance except the $\ell_{1}$ and kkNN score, due to their non-smoothness. Figure 3b shows the ratio of the spectral entropy (Vendi score [16]), which measures sample diversity relative to the observed data. All settings show that nonconformity flows generate samples that are at least as diverse as the target data (ratio $\geq 1$), thus the sampler is not collapsing to a low dimensional subspace even in high dimension. Figure 7 in Appendix D.1 shows multiple realizations from a conformal boundary on 1D functional data to show level set diversity.
Risk Controlling Bands
Symm. Asym. $\Delta\mu$ $\Delta(\mu,\sigma)$
Sample 0.894 0.646 0.828 0.876
Re-conf 0.901 0.900 0.900 0.901
RCPS 0.900 0.901 0.901 0.902
UQNO 0.901 0.902 0.906 0.907
Sample 2.443 2.032 1.145 4.269
Re-conf 2.479 2.948 1.302 4.465
RCPS 2.460 3.487 2.451 5.715
UQNO 2.467 2.982 2.348 5.772
Table 1 shows the marginal coverage ($\alpha=0.1$) and median prediction set width under four 1D functional regression settings Appendix D.1. In the symmetric case, all methods perform the same. For asymmetric and biased data, both mean biased ($\Delta\mu$) and mean+variance biased ($\Delta(\mu,\sigma)$), the nonconformity flow adapts while alternative methods, RCPS [5] and UQNO [29], do not. This is due to our approach conformalizing in the native space of the data, rather than their graphs, thus reflecting its geometric properties. Representative bands are shown in Appendix C.
@@ 4.2 Evaluating Conformal Predictive Distributions
We next compare CPDs to widely used machine-learning ensemble baselines: MC Dropout [17], Deep Ensembles [24], Mean-Variance Networks [33], Conditional Flow Matching [26], Mixture Density Networks [7], and Implicit Quantile Networks [13]. Details of each method provided in Appendix D.2. We consider five structured regression tasks: predict 2D Gaussian processes, inverse model 2D Elliptic PDE solutions, approximate Navier-Stokes solutions ($t=2$), downscale North American precipitation fields, and debiasing a global climate model’s temperature fields (Appendix D.1). We consider both a global $\ell_{2}$ score (CPD-G) and local composite score (CPD-L) (Appendix D.4). We measure the quality of each predictive distribution via Energy Distance (ED) [18], Log Spectral Distance (LSD) [23, 34], and local MMD [6, 3] (Appendix D.6). ED measures overall distribution approximation. LSD measures how energy is distributed over frequency scales, and thus captures pattern and texture similarity, while the local MMD measures local detail fidelity.
@@
Table 2 shows that CPDs achieve competitive ED scores and substantially lower LSD and MMD scores, indicating better alignment with both large-scale structure and fine-scale variability than baseline methods. In each case, the CPD base measure was chosen to be the empirical distribution of the calibration data. Representative samples provided in Section D.7.
Figure 4: UQ spectral decay compared to the targets (GP). CPDs match spectral shape.
@@
Finally, we demonstrate how CPDs enable targeted sampling at arbitrary confidence levels, including extreme-tail and rare-event regimes that are inaccessible to conventional ensemble or conformal approaches. The HURDAT2 dataset records hurricane trajectories from 1850 through 2025 [21]. We train a 1D CNN model to predict the next 12 time steps of the trajectory, given its previous 24 steps and auxiliary information including wind speed and pressure (Appendix D.1). We conformalize the predicted trajectories using a naive $\ell_{2}$ score and conditional geometric trajectory score (CGT) defined in Appendix D.5.
@@ Appendix A Nonconformity velocity field
Among all velocity fields $v(\cdot)$ satisfying 16, we choose the minimum Euclidean norm velocity field
@@ Appendix B Proofs
Appendix C Additional Simulation
C.1 Admissible Representations
@@
C.2 Risk Controlling bands
Figure 8 shows the upper band for each method on each setting, showing how re-conformalization adapts to asymmetry, mean biases, and heteroskedasticity across the domain. Thus, even when the underlying model is heavily misspecified, CPD provides tight, exactly controlled, prediction bands. The four panels correspond to the four columns in Table 1, with the exact data generation defined in Section D.1 under 1D Gaussian processes.
@@ C.2 Risk Controlling bands
Appendix D Simulation Details
D.1 Datasets and processing
@@ D.2 Models
D.3 Vector Nonconformity Scores
@@
D.4 2D Operator Nonconformity Scores
@@
D.5 Trajectory nonconformity scores
@@ D.6 Evaluation Metrics
D.7 Qualitative Samples (Table 2)

### v2

@@ Abstract
Conformal prediction provides a distribution-free framework for uncertainty quantification via prediction sets with exact finite-sample coverage. In low dimensions these sets are easy to interpret, but in high-dimensional or structured output spaces they are difficult to represent and use, which can limit their ability to integrate with downstream tasks such as sampling and probabilistic forecasting. We show that any differentiable nonconformity score induces a deterministic flow on the output space whose trajectories converge to the boundary of the corresponding conformal prediction set. This leads to a computationally efficient, training-free method for sampling conformal boundaries in arbitrary dimensions. Boundary samples can be reconformalized to form pointwise prediction sets with controlled risk and, optionally, repulsed along the boundary to improve geometric coverage. Mixing across confidence levels yields conformal predictive distributions whose quantile regions coincide exactly with conformal prediction sets. We evaluate the approach on PDE inverse problems, precipitation downscaling, climate model debiasing, and hurricane trajectory forecasting.
@@ 1 Introduction
In this work, we propose a constructive alternative based on flows [36]. We show that any differentiable nonconformity score can induce a deterministic dynamical system on $\mathcal{Y}$ whose trajectories globally attract to the conformal set boundary. The resulting flow, called a nonconformity flow, is uniquely determined up to tangential motion by the gradient of the nonconformity score, and, therefore, requires no additional training, auxiliary modeling, or transport mapping. We show that, for almost any initialization, the flow trajectories converge exponentially fast to the conformal boundaries in arbitrary dimensions. Therefore, nonconformity flows provide an exact, computationally efficient, and highly scalable mechanism for sampling conformal prediction set boundaries.
Furthermore, by mixing across confidence levels we can induce calibrated predictive distributions, called conformal predictive distributions (CPDs), for which every conformal prediction set $C_{\alpha}(x)$ is an exact quantile region. CPDs are a scalable Monte Carlo alternative to analytical conformal predictive systems [42] and optimal transport maps [32] for lifting conformal prediction from sets to predictive distributions. We also show how additional tangential dynamics can be used to improve exploration of conformal boundaries when the CPD is overly concentrated on each boundary.
We evaluate nonconformity flows and CPDs on a range of structured regression tasks, including elliptic PDE inverse problems, precipitation downscaling, climate model debiasing, and trajectory forecasting (Section 4). We show that nonconformity flows quickly and reliably converge across a range of data modalities and score functions. CPDs are shown to be competitive with baseline uncertainty quantification (UQ) methods and can act as fully controllable rare-event simulators (Figure 7).
@@ 3 Method
Section 3 develops our method in three stages. First, we introduce nonconformity flows to efficiently sample from $\partial C_{\alpha}(x)=\{y\in\mathcal{Y}:S(x,y)=\tau_{\alpha}\}$ in arbitrary dimensions given any differentiable score (Section 3.1). We then show how to convert these into risk-controlling prediction bands (Section C). Finally, we show how mixing across confidence levels induces a class of calibrated conformal predictive distributions (CPDs), allowing us to transform conformal sets in to fully fledged predictive distributions (Section 3.2).
@@ 3.1 Sampling conformal boundaries
$T_{\varepsilon}(y_{0})=(1/\lambda)\log\big[|S(y(0))-\tau_{\alpha}|/\varepsilon\big]$ (9)
@@ 2. Sample $y_{0}\sim\mu_{x}$
3.3 Tangent Repulsion for Boundary Exploration
Figure 3: Repulsive boundary exploration under a kkNN score (Section E.3).
Because the flow $\Phi_{\alpha}(t,y_{0})$ is everywhere co-linear with $\nabla S$, the induced boundary measures $\{\nu_{x,\alpha}\}_{\alpha\in(0,1)}$ are, in effect, projections of the base measure $\mu_{x}$ onto the level sets $\{\partial C\alpha(x)\}_{\alpha\in(0,1)}$ along integral curves of $\nabla S$. Consequently, not all boundary points are reachable from a given base measure: trajectories are attracted to nearby regions of the boundary and may never explore distant regions (Figure 3). This behavior is partially desirable. For instance, when $\mu_{x}$ is concentrated near the data manifold (e.g., calibration data), then $P_{x}^{\mathrm{CPD}}$ is also close to the manifold. However, this also induces a systematic bias: trajectories tend to accumulate in geometrically flat, nearby regions of the boundary, leading to poor exploration of $\partial C_{\alpha}(x)$.
To mitigate this effect, we introduce a deterministic tangent repulsion velocity field that spreads trajectories along the level set while preserving the score constraint. The minimum-norm controller used to drive trajectories to the score level set $\partial C_{\alpha}(x)$ enforces a prescribed scalar ODE for $S(y(t))$, but does not uniquely specify the velocity field. In particular, any velocity of the form
Algorithm 2 Tangent Repulsion
Input: Boundary points $\{y_{i}(t)\}_{i=1}^{B}$, step size $\delta$
for $k=1$ to $K$ do
  1. Compute $R_{i}(t)=\sum_{j\neq i}(y_{i}-y_{j})/(\|y_{i}-y_{j}\|_{2}^{2})$ and $g_{i}=\nabla S(y_{i})$
  2. Compute $v_{\mathrm{t}}(y_{i})=(I-g_{i}g_{i}^{\top}/\|g_{i}\|^{2})R_{i}$
  3. Step $y_{i}\leftarrow y_{i}+\delta\,v_{\mathrm{t}}(y_{i})$
  4. Correct $y_{i}\leftarrow\lim_{t\rightarrow\infty}\Phi_{\alpha}(t,y_{i})$
end for
Return: Equalized boundary points $\{y_{i}(t)\}_{i=1}^{B}$
$y^{\prime}(t)=v_{\mathrm{n}}(y(t))+v_{\mathrm{t}}(y(t)),\qquad\nabla S(y)^{\top}v_{\mathrm{t}}(y)=0,$ (13)
induces the same evolution of $S(y(t))$ as the minimum-norm flow (Eqn. 7), since $v_{\mathrm{t}}$ lies in the tangent space of $\partial C_{\alpha}(x)$.
Given a batch of boundary points $\{y_{i}(t)\}_{i=1}^{B}$ at time $t$, we define pairwise Euclidean repulsion scores
$R_{i}(t)\;=\;\sum_{j\neq i}\frac{y_{i}(t)-y_{j}(t)}{\|y_{i}(t)-y_{j}(t)\|_{2}^{2}},$ (14)
and project it onto the tangent space of the level set at $y_{i}(t)$ to define a tangential velocity field
$v_{\mathrm{t}}(y_{i}(t))\;=\;\Big(I-\frac{g_{i}g_{i}^{\top}}{\|g_{i}\|_{2}^{2}}\Big)R_{i}(t),$ (15)
where $g_{i}=\nabla S(y_{i})$ is the score gradient. By construction, $S(y)_{i}^{\top}v_{\mathrm{t}}(y_{i})=0$, so $v_{\mathrm{t}}$ does not alter $S(y)$. However, by flowing over $v_{\mathrm{t}}$, we maximize the pairwise Euclidean distance between the initial boundary points. This results in a relatively homogenous sampling of points along $\partial C_{\alpha}(x)$ independent of the base measure $\mu_{0}$ (Algorithm 2). Whether uniform exploration of $\partial C_{\alpha}(x)$ is desirable or not, however, is problem specific, so we consider repulsive exploration as optional.
@@ 4 Numerical Experiments
We evaluate the proposed framework along two axes corresponding (i) convergence and robustness of the nonconformity flow (Section 4.1) and (ii) conformal predictive distributions (CPDs) as calibrated predictive objects (Section 4.2). Definitions of score functions, data generation, preprocessing, model architectures, and evaluation metrics are deferred to Appendices E.1 - E.6.
@@ 4.1 Evaluating Conformal Level Sets
We first evaluate whether the proposed flow (Equation 5) converges to within tolerance ($\varepsilon=10^{-6}$) with $\lambda$ set per trajectory according to the $\varepsilon$-hitting time equation (Equation 3.2). We consider isotropic Gaussian, anisotropic Student-t ($df=3$), 2D Gaussian processes, and 2D GP upsampling targets across a range of dimensions 51005-100 (isotropic Gaussians, anisotropic Student-t) or intrinsic dimension range $\ell=0.01-0.2$ (2D GP tasks) (Appendix E.1). For the vector regression tasks, we train a multilayer perceptron (Appendix E.2) and for the operator tasks, a 2D Fourier Neural Operator (Appendix E.2). All models trained on a single A100 GPU for 25-100 epochs until convergence using the Adam optimizer with learning rate $\gamma=10^{-3}$.
@@
Figure 4a, shows the log score error after 20 integration steps broken out by score function (Appendix E.3E.4) with the black line indicating the target tolerance. Across all settings and tasks the conformal flow sampler converges to within tolerance except the $\ell_{1}$ and kkNN score, due to their non-smoothness. Figure 4b shows the ratio of the spectral entropy (Vendi score [16]), which measures sample diversity relative to the observed data. All settings show that nonconformity flows generate samples that are at least as diverse as the target data (ratio $\geq 1$), thus the sampler is not collapsing to a low dimensional subspace even in high dimension. Figure 8 in Appendix E.1 shows multiple realizations from a conformal boundary on 1D functional data to show level set diversity.
Naive v.s. Repulsive sampling
Table 1: Avg. nearest neighbor distance
Dim. $p=$ $10$ $25$ $50$ $100$
Naive 0.624 0.964 1.124 1.225
Repulse 2.150 4.604 6.296 7.591
[…]
@@ 4.2 Evaluating Conformal Predictive Distributions
We next compare CPDs to widely used machine-learning ensemble baselines: MC Dropout [17], Deep Ensembles [24], Mean-Variance Networks [33], Conditional Flow Matching [26], Mixture Density Networks [7], and Implicit Quantile Networks [13]. Details of each method provided in Appendix E.2. We consider five structured regression tasks: predict 2D Gaussian processes, inverse model 2D Elliptic PDE solutions, approximate Navier-Stokes solutions ($t=2$), downscale North American precipitation fields, and debiasing a global climate model’s temperature fields (Appendix E.1). We consider both a global $\ell_{2}$ score (CPD-G) and local composite score (CPD-L) (Appendix E.4). We measure the quality of each predictive distribution via Energy Distance (ED) [18], Log Spectral Distance (LSD) [23, 34], and local MMD [6, 3] (Appendix E.6). ED measures distribution approximation, LSD measures how energy is distributed over frequency scales, measuring spatial pattern similarity, while the local MMD measures local detail fidelity.
@@
Table 2 shows that CPDs achieve competitive ED scores and substantially lower LSD and MMD scores, indicating better alignment with both large-scale structure and fine-scale variability than baseline methods. In each case, the CPD base measure was chosen to be the empirical distribution of the calibration data. Representative samples provided in Section E.7.
Figure 5: CPDs match spectral shape.
@@
Finally, we demonstrate how CPDs enable targeted sampling at arbitrary confidence levels, including extreme-tail and rare-event regimes that are inaccessible to conventional ensemble or conformal approaches. The HURDAT2 dataset records hurricane trajectories from 1850 through 2025 [21]. We train a 1D CNN model to predict the next 12 time steps of the trajectory, given its previous 24 steps and auxiliary information including wind speed and pressure (Appendix E.1). We conformalize the predicted trajectories using a naive $\ell_{2}$ score and conditional geometric trajectory score (CGT) defined in Appendix E.5.
@@ Appendix A Nonconformity velocity field
Among all velocity fields $v(\cdot)$ satisfying 19, we choose the minimum Euclidean norm velocity field
@@ Appendix B Proofs
Appendix C Risk–controlling Prediction Bands
$P(\ell_{\eta}(X,Y)\leq\delta\bigr)\geq 1-\alpha.$ (24)
Risk Controlling Bands
Symm. Asym. $\Delta\mu$ $\Delta(\mu,\sigma)$
Sample 0.894 0.646 0.828 0.876
Re-conf 0.901 0.900 0.900 0.901
RCPS 0.900 0.901 0.901 0.902
UQNO 0.901 0.902 0.906 0.907
Sample 2.443 2.032 1.145 4.269
Re-conf 2.479 2.948 1.302 4.465
RCPS 2.460 3.487 2.451 5.715
UQNO 2.467 2.982 2.348 5.772
Table 3 shows the marginal coverage ($\alpha=0.1$) and median prediction set width under four 1D functional regression settings Appendix E.1. In the symmetric case, all methods perform the same. For asymmetric and biased data, both mean biased ($\Delta\mu$) and mean+variance biased ($\Delta(\mu,\sigma)$), the nonconformity flow adapts while alternative methods, RCPS [5] and UQNO [29], do not. This is due to our approach conformalizing in the native space of the data, rather than their graphs, thus reflecting its geometric properties. Representative bands are shown in Appendix D.
Appendix D Additional Simulation
D.1 Admissible Representations
@@
D.2 Risk Controlling bands
Figure 9 shows the upper band for each method on each setting, showing how re-conformalization adapts to asymmetry, mean biases, and heteroskedasticity across the domain. Thus, even when the underlying model is heavily misspecified, CPD provides tight, exactly controlled, prediction bands. The four panels correspond to the four columns in Table 3, with the exact data generation defined in Section E.1 under 1D Gaussian processes.
@@ D.2 Risk Controlling bands
Appendix E Simulation Details
E.1 Datasets and processing
@@ E.2 Models
E.3 Vector Nonconformity Scores
@@
E.4 2D Operator Nonconformity Scores
@@
E.5 Trajectory nonconformity scores
@@ E.6 Evaluation Metrics
E.7 Qualitative Samples (Table 2)